Ellipsis, Tense and Questions

Hans Kamp (editor)

DYANA-2

Dynamic Interpretation of Natural Language

esprit Basic Research Project 6852 Deliverable R2.2.B September 1994

Ellipsis, Tense and Questions

Hans Kamp Universität Stuttgart

(editor)

authors: Helle Frisak Sem Dorit Abusch Irene Heim

Jonathan Ginzburg

Henk Zeevat Bernhard Schwarz Regine Eckardt

Uwe Reyle

DYANA-2

Dynamic Interpretation of Natural Language

esprit Basic Research Project 6852 Deliverable R2.2.B September 1994

esprit Basic Research Project 6852 dyana-2

Dynamic Interpretation of Natural Language

Universiteit van Amsterdam Institute for Logic, Language and Computation (ILLC)

University of Edinburgh Centre for Cognitive Science (ECCS)

Universität München Centrum für Informations- und Sprachforschung (CIS)

Universitetet i Oslo Department of Linguistics and Philosophy (ILF)

Universität Stuttgart Institut für Maschinelle Sprachverarbeitung (IMS)

Universität Tübingen Seminar für Sprachwissenschaft (SfS)

Universiteit Utrecht Research Institute for Language and Speech (OTS)

For copies of reports, updates on project activities and other dyana-related information, contact:

The dyana-2 Project Administrator ILLC/Department of Philosophy University of Amsterdam Nieuwe Doelenstraat 15 NL-1012 CP Amsterdam

\begin{matrix}c \\ ◯ \end{matrix} 1994, The Individual Authors

No part of this document may be reproduced or transmitted in any form, or by any means, electronic or mechanical, including photocopy, recording, or any information storage and retrieval system, without permission from the copyright owner.

Contents

Task 2.2, subtask 4: VP-ellipsis VP-ellipsis and DRT Helle Frisak Sem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Task 2.2, subtask 5: Tense in Intensional Contexts Sequence of Tense Revisited: Two Semantic Accounts of Tense in Intensional Contexts Dorit Abusch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

Comments on Abusch’s Theory of Tense Irene Heim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

Task 2.2, subtask 7: Questions Resolving Questions Jonathan Ginzburg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171

Task 2.2, subtask 7: Questions Applying an Exhaustivity Operator in Update Semantics Henk Zeevat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

Additional Contribution A Note on Uniqueness with Questions Bernhard Schwarz. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269

Additional Contribution Adverbs in Focus Regine Eckardt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289

Additional Contribution Monotonic Disambiguation and Plural Pronoun Resolution Uwe Reyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333

Task 2.2, subtask 4 VP-ellipsis

VP-ellipsis and DRT

Helle Frisak Sem

(Universitetet i Oslo)

Task 2.2, subtask 5

Tense in Intensional Contexts

Sequence of Tense Revisited: Two Semantic Accounts of Tense in

Intensional Contexts

Dorit Abusch

(Universität Stuttgart)

Comments

Comments on Abusch’s Theory of Tense

Irene Heim

(Massachusetts Institute of Technology)

Task 2.2, subtask 7 Questions

Resolving Questions

Jonathan Ginzburg

(University of Edinburgh)

Applying an Exhaustivity Operator in Update Semantics

Henk Zeevat

(Universiteit van Amsterdam)

Additional Contribution

A Note on Uniqueness with Questions

Bernhard Schwarz

(Universität Tübingen)

Additional Contribution

Adverbs in Focus

Regine Eckardt

(Universität Tübingen)

Additional Contribution

Monotonic Disambiguation and Plural Pronoun Resolution

Uwe Reyle

(Universität Stuttgart)